# In Times of Crisis: An Exploratory Study of Media and Political Discourse on YouTube During the 2024 French Elections

### v1

@@ Abstract
YouTube has emerged as a major platform for political communication and news dissemination, particularly during high-stakes electoral periods. In the context of the 2024 European Parliament and French legislative elections, this study investigates how political actors and news media used YouTube to shape public discourse. We analyze over 100,000 video transcripts and metadata from 74 French YouTube channels operated by national news outlets, local media, and political figures. To identify the key themes emphasized during the campaign period, we applied a semi-automated method that combined large language models with clustering and manual review. The results reveal distinct thematic patterns across the political spectrum and media types, with right-leaning news outlets focusing on topics like immigration, while left-leaning emphasized protest and media freedom. Themes generating the most audience engagement, measured by comment-to-view ratios, were most often the most polarizing ones. In contrast, less polarizing themes such as video games and nature showed higher approval, reflected in like-to-view ratios. We also observed a general tendency across all media types to portray political figures in neutral or critical terms rather than favorable ones.
@@ 1 Introduction
YouTube has become an increasingly important platform for news consumption, particularly among younger audiences who turn away from traditional media. According to the 2024 Reuters Institute Digital News Report, there is a clear global shift toward video-based content, with platforms like YouTube gaining traction as news and political commentary sources (reuters). Its accessibility, free use, and algorithm-driven personalization make it especially attractive to users seeking real-time information and a wider range of perspectives than those typically offered by mainstream broadcast outlets. With over two billion active users globally each month, YouTube serves not only as a major social media platform but also as a key source of news and information. Moreover, YouTube offers political parties and politicians a powerful platform to communicate directly with the public, bypassing the editorial gatekeeping of traditional broadcast media. This autonomy allows them to craft and control their messaging, tailor content to specific audiences, and respond rapidly to unfolding events. As a result, YouTube has become a strategic tool for promoting campaign narratives and shaping political discourse on its terms (30channels; ukpolit). YouTube’s popularity, accessibility, rich metadata, and mix of political and media actors make it a powerful platform for analyzing and comparing evolving patterns of political communication and news media discourse.
While previous research has examined political discourse during elections across various social media platforms and countries (oliveira2018politicians; wu2021cross; covidfat; 10.1145/3665450; 10.1145/3543507.3583875; caro), studies focusing specifically on political and news media discourse on YouTube during election periods remain scarce. In this context, the French 2024 elections present a unique and timely case study, shaped by political crisis and the rise of the far right. The European Parliament election, held on 9 June 2024, triggered a major political upheaval following the far-right Rassemblement National’s best-ever result (brookings), surpassing President Macron’s Renaissance party. In response, Macron dissolved the National Assembly and called snap legislative elections. This period of political instability offers a valuable opportunity to examine how discourse unfolded across French news media and political communication on YouTube.
@@
For this study, we analyzed YouTube channels operated by national and local news media, political parties, and politicians who participated in the 2024 European and French legislative elections. To capture both the buildup and immediate aftermath of these events, we analyzed content published between 1 March 2023 and 15 July 2024. Using the YouTube API to collect channel and video metadata, and the Transcript API to extract video transcripts, we compiled a dataset of over 100,000 video transcripts and related metadata from 74 selected channels.
A central methodological challenge in analyzing political discourse during the 2024 French elections is identifying which themes to capture and whether existing frameworks are suitable for this context. In this study, we chose not to rely on predefined codebooks such as those from the Comparative Agenda Project (CAP) (Baumgartner-2006) or the Comparative Manifesto Project (CMP) (CMP), despite their wide use in political content analysis. These frameworks are designed for long-term agenda tracking and may miss the breadth and specificity of themes that surfaced in news coverage and political messaging during the 2024 election period. Predefined categories also risk overlooking emergent or context-specific issues that were particularly salient at the time.
To extract themes from the transcripts, we employed a multi-step, semi-automated approach. First, we used a large language model (Mistral’s NeMo) to generate topic labels for each transcript (nemo). The number of topics requested was proportionate to the transcript’s length, ensuring appropriate granularity. These topics were then vectorized and clustered using the k-means algorithm to identify recurring themes across the dataset. The resulting clusters were manually reviewed and refined to merge semantically similar themes and improve interpretability.
@@ 2 Data Collection
We compiled a list of parties and politicians participating in the 2024 European and French legislative elections (europ_el; leg_el) and manually searched for their official YouTube channels. For the European elections, we focused on party leaders and representatives whose parties secured seats in the European Parliament. For the legislative elections, we selected prominent participants, including major party leaders and high-profile candidates from across the political spectrum. In both cases, we began by checking official websites for links to YouTube channels; if none were available, we conducted manual searches using the individuals’ names directly on YouTube. A channel was included in our dataset if it had more than 10 videos posted during the relevant campaign period. In total, we collected YouTube channels for 17 politicians and 10 parties.
@@
To identify YouTube channels of national news media, we began with the Media Bias/Fact Check list, which evaluates sources based on their editorial content and reporting practices (mediabias). As this list was incomplete for French outlets, we supplemented it with sources recognized by the Conseil supérieur de l’audiovisuel (CSA) France’s former regulatory authority for audiovisual media  (CSA). We included only channels that consistently published informational content, not solely entertainment, over a three-month period and had at least 10,000 subscribers. In total, we identified channels of 36 national news sources that met our requirements.
@@
For local news channels, we used the list provided by the Alliance pour les Chiffres de la Presse et des Médias  (ACPM), which tracks circulation and audience data for regional print and online media. The list included 62 regional outlets, of which we located YouTube channels for 43. We selected those with over 5,000 subscribers and at least some informational content published over a six-month period. In total, 15 regional channels met our criteria and were included for the next step of data collection.
@@
We identified the political orientations of politicians and their affiliated parties using Légifrance, the official French government website for legal information (Legifarnce). Determining the political orientation of news channels was challenging due to limited official information and the evolving nature of media bias. To support our classifications, we relied on references from Media Bias/Fact Check and the European media monitoring platform Eurotopics (eurotop).
@@
After finalizing the selection of YouTube channels, we utilized the YouTube API to gather detailed channel metadata, including the number of videos, subscribers, and total likes, to assess the scale and engagement of each channel (youtube-api). Using the same API, we retrieved video IDs for all videos published by these channels within the timeframe from March 1, 2023, to July 15, 2024 (to monitor time periods with and without elections). Additionally, we collected the metadata of each video: title, description, view count, like count, comment count, and publication date. Metadata was collected in September, about two months after the last video in the dataset was published.
@@
Next, using the video IDs and the YouTube Transcript API, we retrieved video transcripts (transcript-api). Manually created transcripts were prioritized over automatically generated ones, and they were downloaded when available. In cases they were unavailable, auto-generated transcripts were collected. For videos without any available transcripts, no further attempts were made to obtain them. Since the data was publicly available and did not involve private individuals or sensitive information, individual consent was not required under current research ethics guidelines. From party and politician channels (for simplicity, we refer to them as political channels throughout the paper), we collected 4,678 transcripts out of 5,435 videos (Political dataset). For national news media channels, we obtained 100,582 transcripts from a total of 150,843 videos (News dataset), and for regional news media, we collected 4,657 transcripts from 8,873 videos (Local dataset). The considered channels and their statistics are shown in Appendix Tables 14, 15 and 16. Some transcripts were unavailable, either because the videos contained no spoken language or because the channels disabled auto-captioning and did not provide their own. Even though some transcripts are missing, the dataset still includes many videos from each main political group, making meaningful comparisons possible.
@@ 3 Methodology for Themes Extraction
We started our theme extraction with a classic topic modeling approach using BERTopic (bertopic). To improve granularity, we split transcripts into 500-word segments and ran the algorithm multiple times. However, the results were unsatisfactory: all topics related to French politics were collapsed into a single category, and many of the extracted topics were difficult to interpret.
In parallel, recent studies have shown that large language models (LLMs) can be effectively employed for topic extraction (mu2024largelanguagemodelsoffer; doi-etal-2024-topic; Invernici_2025; CELIKTEN2025113219). To perform topic labeling on our transcripts, we selected a suitable large language model (LLM) based on four main criteria. The model needed strong performance on labeling tasks to ensure accuracy, and training on French-language data to match our transcripts. We prioritized open-source models for flexibility and transparency, and considered cost-efficiency, given the computational demands of large-scale processing. Considering all these aspects, we selected NeMo (nemo), an LLM developed by Mistral.
@@
After obtaining topics, we used Mistral embeddings to represent the topics in a high-dimensional space and applied k-means clustering to group similar topics (ms_embedding). To determine the number of clusters, we tested the elbow method and silhouette analysis. Although these methods produced reasonable results, they tended to merge distinct topics that we deemed should be kept separate (such as Macron’s Governance and Zemmour). Since the dataset can be meaningfully partitioned at different levels of granularity, we opted for a larger number of clusters to preserve fine distinctions and allow for flexibility for later aggregation into broader themes.
@@ 4.1 Analysis of Theme Distribution
We also included an analysis of local news sources, as prior studies indicate that people tend to place greater trust in local news providers compared to national or official outlets (local). Given that the dataset includes only 4,000 videos from local sources, we chose not to divide them by time periods or political orientation.
@@ 4.4 Thematic Landscape of YouTube Channels
To explore whether YouTube channels of news outlets and political figures tended to align along editorial lines, we visualized their thematic representations. We selected channels that have twenty or more videos in our dataset. For each channel, we constructed a theme vector by counting how often each theme appeared across its videos, counting all unique themes per video. These raw counts were then normalized to produce a probability distribution over themes for each channel. We then applied the dimensionality reduction technique t-SNE to project the theme vectors into a two-dimensional space for visualization, as it is well-suited for preserving local structure and revealing clustering patterns in high-dimensional data (tsne). The appendix provides details on the validation of the visualization results.
@@ 4.5 Stance Detection
To select transcripts for stance detection, we used the RapidFuzz Python library (fuzz), which enables fuzzy matching, useful for handling misspellings and transcription errors common in auto-generated YouTube transcripts. Due to a low number of relevant transcripts for Ciotti (fewer than 50), we excluded him from the stance detection analysis.
@@ 5 Related Work
Studies about YouTube. YouTube has become a central platform for political communication, with US users reportedly watching 100 times more candidate-related videos on YouTube than on CNN, C-SPAN, MSNBC, and Fox News combined (GOOGLE). As a result, political video content has drawn increasing scholarly attention. Prior works have addressed ad classification challenges using multimodal approaches to handle tasks such as tone detection and distinguishing political ads from other online content (imabalnce_data; video_classification).
@@
In parallel, a growing body of work has examined harmful and extremist content on YouTube. rad_patways analyzed over 330,000 videos and found that users often migrate toward more radical content over time.  rec_system explored the role of YouTube’s recommendation algorithm in amplifying far-right content, raising concerns about algorithmic bias and radicalization pathways.
@@
YouTube’s comment sections have also become a focus of research. jiang2019bias examine the dynamics of comment moderation on YouTube, focusing on the perceived political bias in content removal. Their findings suggest that higher moderation rates are more strongly linked to ideological extremism and misinformation than to political orientation alone, challenging common assumptions about partisan censorship on the platform. doi:10.1177/21582440221094587 proposed a user activity-based model to assess political polarization on South Korean YouTube channels, analyzing over 11 million comments from 600,000 users. Their study revealed a high degree of polarization, with only 8% of users identified as neutral. wu2021cross examined 134 million YouTube comments and found that while cross-partisan engagement exists, it is asymmetrical; conservatives are significantly more likely to engage with liberal content than liberals with conservative content.
@@
Finally,  10.1145/3614419.3644023 presented the first large-scale comparison of YouTube Shorts and regular videos, analyzing over 17 million videos across 70,000 channels. Their findings show that Shorts are popular among new creators and are mainly used for entertainment, while regular videos remain dominant in educational and political domains.
@@
Studies about online news media. Many studies have examined online news, focusing on the dynamics of information flows  (RichardFletcher2020; doi:10.1126/sciadv.aau4586; Ognyanova2020; efstratiou2022adherence; papadamou2022just; mosleh2022measuring; Horne_Gruppi; 10.1145/3757540). Complementing these works, 10646651 developed an automated method to identify and track misleading narratives across low-credibility news websites. Their analysis revealed that lesser-known outlets play a significant role in spreading misinformation and demonstrated how automated systems can support an early detection of emerging false narratives. chouaki addressed limitations in platform access and introduced a data donation approach that captures users’ interactions with news content. Their findings suggest that misinformation exposure is often driven more by users’ own content choices than by algorithms.

### v2

@@ Abstract
YouTube has emerged as a major platform for political communication and news dissemination, particularly during high-stakes electoral periods. In the context of the 2024 European Parliament and French legislative elections, this study investigates how political actors and news media used YouTube to shape public discourse. We analyze over 100,000 video transcripts and metadata from 73 French YouTube channels operated by national news outlets, local media, and political figures. To identify the key themes emphasized during the campaign period, we applied a semi-automated method that combined large language models with clustering and manual review. The results reveal distinct thematic patterns across the political spectrum and media types, with right-leaning news outlets focusing on topics like immigration, while left-leaning emphasized protest and media freedom. Themes generating the most audience engagement, measured by comment-to-view ratios, were most often the most polarizing ones. In contrast, less polarizing themes such as video games and nature showed higher approval, reflected in like-to-view ratios. We also observed a general tendency across all media types to portray political figures in neutral or critical terms rather than favorable ones.
@@ 1 Introduction
YouTube has become an increasingly important platform for news consumption, particularly among younger audiences who turn away from traditional media. According to the 2024 Reuters Institute Digital News Report, there is a clear global shift toward video-based content, with platforms like YouTube gaining traction as news and political commentary sources (Reuters 2024). Its accessibility, free use, and algorithm-driven personalization make it especially attractive to users seeking real-time information and a wider range of perspectives than those typically offered by mainstream broadcast outlets. With over two billion active users globally each month, YouTube serves not only as a major social media platform but also as a key source of news and information. Moreover, YouTube offers political parties and politicians a powerful platform to communicate directly with the public, bypassing the editorial gatekeeping of traditional broadcast media. This autonomy allows them to craft and control their messaging, tailor content to specific audiences, and respond rapidly to unfolding events. As a result, YouTube has become a strategic tool for promoting campaign narratives and shaping political discourse on its terms (Tubics 2024; Glick 2024). YouTube’s popularity, accessibility, rich metadata, and mix of political and media actors make it a powerful platform for analyzing and comparing evolving patterns of political communication and news media discourse.
While previous research has examined political discourse during elections across various social media platforms and countries (Oliveira et al. 2018; Wu and Resnick 2021; Shi et al. 2022; Noor et al. 2024; Sosnovik et al. 2023; Violot et al. 2026), studies focusing specifically on political and news media discourse on YouTube during election periods remain scarce. In this context, the French 2024 elections present a unique and timely case study, shaped by political crisis and the rise of the far right. The European Parliament election, held on 9 June 2024, triggered a major political upheaval following the far-right Rassemblement National’s best-ever result (Varma 2024), surpassing President Macron’s Renaissance party. In response, Macron dissolved the National Assembly and called snap legislative elections. This period of political instability offers a valuable opportunity to examine how discourse unfolded across French news media and political communication on YouTube.
@@
For this study, we analyzed YouTube channels operated by national and local news media, political parties, and politicians who participated in the 2024 European and French legislative elections. To capture both the buildup and immediate aftermath of these events, we analyzed content published between 1 March 2023 and 15 July 2024. Using the YouTube API to collect channel and video metadata, and the Transcript API to extract video transcripts, we compiled a dataset of over 100,000 video transcripts and related metadata from 73 selected channels.
A central methodological challenge in analyzing political discourse during the 2024 French elections is identifying which themes to capture and whether existing frameworks are suitable for this context. In this study, we chose not to rely on predefined codebooks such as those from the Comparative Agenda Project (CAP) (Baumgartner et al. 2006) or the Comparative Manifesto Project (CMP) (Volkens et al. 2020), despite their wide use in political content analysis. These frameworks are designed for long-term agenda tracking and may miss the breadth and specificity of themes that surfaced in news coverage and political messaging during the 2024 election period. Predefined categories also risk overlooking emergent or context-specific issues that were particularly salient at the time.
To extract themes from the transcripts, we employed a multi-step, semi-automated approach. First, we used a large language model (Mistral’s NeMo) to generate topic labels for each transcript (Mistral AI team 2024b). The number of topics requested was proportionate to the transcript’s length, ensuring appropriate granularity. These topics were then vectorized and clustered using the k-means algorithm to identify recurring themes across the dataset. The resulting clusters were manually reviewed and refined to merge semantically similar themes and improve interpretability.
@@ 2 Data Collection
We compiled a list of parties and politicians participating in the 2024 European and French legislative elections (Wikipedia 2024a, b) and manually searched for their official YouTube channels. For the European elections, we focused on party leaders and representatives whose parties secured seats in the European Parliament. For the legislative elections, we selected prominent participants, including major party leaders and high-profile candidates from across the political spectrum. In both cases, we began by checking official websites for links to YouTube channels; if none were available, we conducted manual searches using the individuals’ names directly on YouTube. A channel was included in our dataset if it had more than 10 videos posted during the relevant campaign period. In total, we collected YouTube channels for 17 politicians and 10 parties.
@@
To identify YouTube channels of national news media, we began with the Media Bias/Fact Check list, which evaluates sources based on their editorial content and reporting practices (Zandt 2015). As this list was incomplete for French outlets, we supplemented it with sources recognized by the Conseil supérieur de l’audiovisuel (CSA) France’s former regulatory authority for audiovisual media  (Conseil supérieur de l’audiovisuel 2024). We included only channels that consistently published informational content, not solely entertainment, over a three-month period and had at least 10,000 subscribers. In total, we identified channels of 36 national news sources that met our requirements.
@@
For local news channels, we used the list provided by the Alliance pour les Chiffres de la Presse et des Médias  (ACPM 2024), which tracks circulation and audience data for regional print and online media. The list included 62 regional outlets, of which we located YouTube channels for 43. We selected those with over 5,000 subscribers and at least some informational content published over a six-month period. In total, 15 regional channels met our criteria and were included for the next step of data collection.
@@
We identified the political orientations of politicians and their affiliated parties using Légifrance, the official French government website for legal information (Legifrance 2023). Determining the political orientation of news channels was challenging due to limited official information and the evolving nature of media bias. To support our classifications, we relied on references from Media Bias/Fact Check and the European media monitoring platform Eurotopics (Eurotopics 2023).
@@
After finalizing the selection of YouTube channels, we utilized the YouTube API to gather detailed channel metadata, including the number of videos, subscribers, and total likes, to assess the scale and engagement of each channel (Google 2023). Using the same API, we retrieved video IDs for all videos published by these channels within the timeframe from March 1, 2023, to July 15, 2024 (to monitor time periods with and without elections). Additionally, we collected the metadata of each video: title, description, view count, like count, comment count, and publication date. Metadata was collected in September, about two months after the last video in the dataset was published.
@@
Next, using the video IDs and the YouTube Transcript API, we retrieved video transcripts (PyPi 2023). Manually created transcripts were prioritized over automatically generated ones, and they were downloaded when available. In cases they were unavailable, auto-generated transcripts were collected. For videos without any available transcripts, no further attempts were made to obtain them. Since the data was publicly available and did not involve private individuals or sensitive information, individual consent was not required under current research ethics guidelines. From party and politician channels (for simplicity, we refer to them as political channels throughout the paper), we collected 4,678 transcripts out of 5,435 videos (Political dataset). For national news media channels, we obtained 100,582 transcripts from a total of 150,843 videos (News dataset), and for regional news media, we collected 4,657 transcripts from 8,873 videos (Local dataset). The considered channels and their statistics are shown in Appendix Tables 14, 15 and 16. Some transcripts were unavailable, either because the videos contained no spoken language or because the channels disabled auto-captioning and did not provide their own. Even though some transcripts are missing, the dataset still includes many videos from each main political group, making meaningful comparisons possible.
@@ 3 Methodology for Themes Extraction
We started our theme extraction with a classic topic modeling approach using BERTopic (Grootendorst 2022). To improve granularity, we split transcripts into 500-word segments and ran the algorithm multiple times. However, the results were unsatisfactory: all topics related to French politics were collapsed into a single category, and many of the extracted topics were difficult to interpret.
In parallel, recent studies have shown that large language models (LLMs) can be effectively employed for topic extraction (Mu et al. 2024; Doi et al. 2024; Invernici et al. 2025; Çelikten and Onan 2025). To perform topic labeling on our transcripts, we selected a suitable large language model (LLM) based on four main criteria. The model needed strong performance on labeling tasks to ensure accuracy, and training on French-language data to match our transcripts. We prioritized open-source models for flexibility and transparency, and considered cost-efficiency, given the computational demands of large-scale processing. Considering all these aspects, we selected NeMo (Mistral AI team 2024b), an LLM developed by Mistral.
@@
After obtaining topics, we used Mistral embeddings to represent the topics in a high-dimensional space and applied k-means clustering to group similar topics (Mistral AI team 2024a). To determine the number of clusters, we tested the elbow method and silhouette analysis. Although these methods produced reasonable results, they tended to merge distinct topics that we deemed should be kept separate (such as Macron’s Governance and Zemmour). Since the dataset can be meaningfully partitioned at different levels of granularity, we opted for a larger number of clusters to preserve fine distinctions and allow for flexibility for later aggregation into broader themes.
@@ 4.1 Analysis of Theme Distribution
We also included an analysis of local news sources, as prior studies indicate that people tend to place greater trust in local news providers compared to national or official outlets (Achouch 2024). Given that the dataset includes only 4,000 videos from local sources, we chose not to divide them by time periods or political orientation.
@@ 4.4 Thematic Landscape of YouTube Channels
To explore whether YouTube channels of news outlets and political figures tended to align along editorial lines, we visualized their thematic representations. We selected channels that have twenty or more videos in our dataset. For each channel, we constructed a theme vector by counting how often each theme appeared across its videos, counting all unique themes per video. These raw counts were then normalized to produce a probability distribution over themes for each channel. We then applied the dimensionality reduction technique t-SNE to project the theme vectors into a two-dimensional space for visualization, as it is well-suited for preserving local structure and revealing clustering patterns in high-dimensional data (van der Maaten and Hinton 2008). The appendix provides details on the validation of the visualization results.
@@ 4.5 Stance Detection
To select transcripts for stance detection, we used the RapidFuzz Python library (PyPi 2024), which enables fuzzy matching, useful for handling misspellings and transcription errors common in auto-generated YouTube transcripts. Due to a low number of relevant transcripts for Ciotti (fewer than 50), we excluded him from the stance detection analysis.
@@ 5 Related Work
Studies about YouTube. YouTube has become a central platform for political communication, with US users reportedly watching 100 times more candidate-related videos on YouTube than on CNN, C-SPAN, MSNBC, and Fox News combined (Stanford 2016). As a result, political video content has drawn increasing scholarly attention. Prior works have addressed ad classification challenges using multimodal approaches to handle tasks such as tone detection and distinguishing political ads from other online content (Baskota 2019; Banerjee 2017).
@@
In parallel, a growing body of work has examined harmful and extremist content on YouTube. Ribeiro et al. (2020) analyzed over 330,000 videos and found that users often migrate toward more radical content over time.  O’Callaghan et al. (2015) explored the role of YouTube’s recommendation algorithm in amplifying far-right content, raising concerns about algorithmic bias and radicalization pathways.
@@
YouTube’s comment sections have also become a focus of research. Jiang et al. (2019) examine the dynamics of comment moderation on YouTube, focusing on the perceived political bias in content removal. Their findings suggest that higher moderation rates are more strongly linked to ideological extremism and misinformation than to political orientation alone, challenging common assumptions about partisan censorship on the platform. Tran et al. (2022) proposed a user activity-based model to assess political polarization on South Korean YouTube channels, analyzing over 11 million comments from 600,000 users. Their study revealed a high degree of polarization, with only 8% of users identified as neutral. Wu and Resnick (2021) examined 134 million YouTube comments and found that while cross-partisan engagement exists, it is asymmetrical; conservatives are significantly more likely to engage with liberal content than liberals with conservative content.
@@
Finally,  Violot et al. (2024) presented the first large-scale comparison of YouTube Shorts and regular videos, analyzing over 17 million videos across 70,000 channels. Their findings show that Shorts are popular among new creators and are mainly used for entertainment, while regular videos remain dominant in educational and political domains.
@@
Studies about online news media. Many studies have examined online news, focusing on the dynamics of information flows  (Fletcher and Nielsen 2018; Guess et al. 2019; Ognyanova et al. 2020; Efstratiou and De Cristofaro 2022; Papadamou et al. 2022; Mosleh and Rand 2022; Horne et al. 2022; Amieur et al. 2025). Complementing these works, Hanley et al. (2024) developed an automated method to identify and track misleading narratives across low-credibility news websites. Their analysis revealed that lesser-known outlets play a significant role in spreading misinformation and demonstrated how automated systems can support an early detection of emerging false narratives. Chouaki et al. (2024) addressed limitations in platform access and introduced a data donation approach that captures users’ interactions with news content. Their findings suggest that misinformation exposure is often driven more by users’ own content choices than by algorithms.